OpenStudy (anonymous):

HELP! MEDAL! Find the numbers such that four less than half the number is at least five and at most ten. Graph: 1 | 74-95 mi/hr 2 | 96-110 mi/hr 3 | 111 - 130 mi/hr 4 | 131 - 155 mi/hr 5 | 156+ mi/hr @freckles

OpenStudy (freckles):

let us call the numbers x

OpenStudy (freckles):

we have 4 less than half x IS AT LEAST five and AT MOST 10 these parts I put in caps are the inequalities

OpenStudy (freckles):

example: u is AT LEAST 2 means u can be 2 or more so the inequality for "u is at least 2" is u>=2 example: u at AT MOST 2 means u can be 2 or less so the inequality for "u is at most 2" is u<=2

OpenStudy (freckles):

examples for other parts of the sentences: 1) half of 6 is represented as 6/2 2) half of n is represented as x/2 3) "n less than m" means m-n

OpenStudy (freckles):

so you have two inequalities to set up: 4 less than half x is at least 5 and 4 less than half x is at most 10

OpenStudy (freckles):

do you think you translate both expressions there

OpenStudy (anonymous):

3 < x/2 > 5

OpenStudy (freckles):

that inequalitiy doesn't seem right since 3< ...> 5 is not correct this is like saying 3 is less than 5 and 3 is greater than 5

OpenStudy (anonymous):

I dont know then

OpenStudy (freckles):

4 less than half x is at least 5 translates to?

OpenStudy (freckles):

half x is the same thing as saying x/2 4 less than half x is the same thing as saying x/2-4

OpenStudy (freckles):

is at least 5 means >=5

OpenStudy (freckles):

you have x/2-4>=5 and the other inequality is x/2-4<=10

OpenStudy (freckles):

you could write this as : \[10 \ge \frac{x}{2}-4 \ge 5\]

OpenStudy (anonymous):

I see how would we figure our the question

OpenStudy (freckles):

by solving the inequality

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

\[28 \ge x \ge 18\]

OpenStudy (freckles):

sounds great

OpenStudy (anonymous):

Do i need to graph?

OpenStudy (freckles):

i'm not really sure i don't see those instructions

OpenStudy (freckles):

i just see find the numbers

OpenStudy (freckles):

we found the numbers

OpenStudy (freckles):

all the numbers between 18 and 28 including both 18 and 28.

OpenStudy (anonymous):

ok thank you again